OpenStudy (tennistar):

What is the simplified form of the expression 5the square root of v to the eighth power. - the square root of v cubed. - the square root of the quantity 16 times v to the eighth power. + 3the square root of v cubed.? Answer 2vthe square root of v to the fifth power. 9v6the square root of v. v4 + 2vthe square root of v. 9v4the square root of v cubed.

OpenStudy (tennistar):

What is the exact value of the expression 11the square root of 8. + 6the square root of 12. - 5the square root of 2.

OpenStudy (kropot72):

\[5\sqrt{v ^{8}}-\sqrt{v ^{3}}-\sqrt{16v ^{8}}+3\sqrt{v ^{3}}\] The first step is to remove 16 from under the radical sign. What do you get when you do this?

OpenStudy (tennistar):

Canyou help me answer the seocnd question

OpenStudy (tennistar):

I really need help and fast

OpenStudy (tennistar):

like in two minutes

OpenStudy (tennistar):

please

OpenStudy (tennistar):

Please help me

OpenStudy (kropot72):

\[11\sqrt{8}+6\sqrt{12}-5\sqrt{2}\] First step is to get all three terms as products of the square root of 2. Can you do that?

OpenStudy (ganpat):

option c , is the answer...

OpenStudy (ganpat):

\[5v4−v√v−4v4+3v√v = v4+2v√v \] just simplify it..